# How Books Tell a History of Statistics in Portugal: Works of Foreigners, $\boldsymbol{Estrangeirados}$, and Others

### v1

@@ 1 Introduction
The teaching of Statistics descending from the links to Political Economy persisted at the Faculty of Law of the University of Coimbra, giving rise to the publication of various texts (Gouveia et al, 1913; Marnoco e Sousa(10)(10)(10)  The incomplete copy from FDUL did not allow us to access some parts of the book; it should be noted that it has a section on “the sceptics of Statistics”, a subject that is currently widely published. In Chapter III (Statistical Laws), sections 59 and 60 are respectively “Law of large numbers” and “Law of small numbers”. Poisson, wanted his result on the convergence of a succession of binomials to Poisson to be known as the law of small numbers, renamed “law of large numbers” Jacques Bernoulli’s “golden theorem” on the convergence of arithmetic means towards the expectation. Only the expression “law of large numbers” persisted., 1913; Silva and Rocha, 1919; Rocha and Monteiro, 1923; Pereira, 1929).
On the other hand, at the Anthropology Institute of the University of Coimbra, Eugénio Tamagnini had an innovative perspective on the contribution of Statistics in Biometrics research. In Revista da Faculdade de Ciências da Universidade de Coimbra, Tamagnini (1937) published an innovative work in Portugal on analysis of variance. Tamagnini also hired W. L. Stevens, a disciple of Ronald Fisher. Xavier da Cunha (1992) [36] comments that “Tamagnini’s interest in the statistical treatment of data obtained from various populations and in comparing them led him to hire a foreign professor of statistics, Professor W. L. Stevens, who collaborated with him on blood groups and also organised free courses in statistics for students of the exact and natural sciences. Stevens had been a brilliant disciple and collaborator of Dr. Fisher, having stayed in Coimbra ‘making us familiar with statistical methods and their most convenient application’, in Tamagnini’s words. This initiative was undoubtedly very useful.
@@
Fonseca (2011) [38] states that “It is in this context of the necessity to improve the statistical treatment of anthropometric, physiological and demographic data, that one understands the arrival of W. L. Stevens, assigned ‘as a statistician, to guide the application work of modern statistical methods to Biological Sciences and organise initiation courses for professors and students of the same methods’. The ‘Elementary Course of Modern Statistical Methods Applicable to Scientific Investigation’, headed by professor W. L. Stevens, functioned between 1942 and 1944, and is linked to a set of published studies in the series ‘Questões de Método’ [‘Methodogical Issues’]. […] As for the years W. L. Stevens spent in Coimbra, the data which allows us to inquire about his teaching activity is frankly scarce. We know that the course he led functioned between 1942 and 1944, but the apparent absence of Stevens’ curriculum vitae in the archives of the University of Coimbra, and, moreover, the inexistence of annuals of the same institution for the period between 1943 and 1947 (encompassing, thus, the years during which the course led by W. L. Stevens was lectured), conditioned our investigation. What was the programme of the course? Who attended it? What bibliography was recommended? Why was it extinct after only two years? These are just some of the questions that (at least) for now, will remain without an answer.”.
Between 1942 and 1945 Stevens published a remarkable set of works which he categorised as Questions of Method — namely on the $\chi^{2}$ significance test, discriminant analysis, statistical distributions, parameters estimation —, disseminating innovations in mathematical statistics by Karl Pearson, Ronald Fisher, Egon Pearson, Jerszy Neyman and their schools.
Stevens also published, with Aureliano Quintanilha and Hugo Ribeiro (1942), an interesting article on the application of Probability Calculus in the investigation of non sexual reproduction by plasmogamy followed by karyogamy in the higher fungi basidiomycetes, in the Gazeta de Matemática (both in the index and in the article the initials W and L are interchanged).
@@
  • Manuel Zaluar Nunes (1907-1967), who started as an assistant at ISCEF, before going to study in Paris as a fellow of the Instituto de Alta Cultura, published in 1933 with Mário Santos a concise exposition of Elementos de Cálculo das Probabilidades e de Estatística Matemática [Elements of Probability Calculus and Mathematical Statistics]: written in accordance with the program of the 2nd chair of the Instituto Superior de Ciências Económicas e Financeiras. Founder of the Gazeta de Matemática, Portugaliae Mathematica, and the Sociedade Portuguesa de Matemática [Portuguese Mathematical Society], he was an Assistant Professor at FCUL and later a Full Professor at ISA, a position from which he was dismissed in 1947 for political reasons, having to migrate first to France, where he was a research assistant at the Centre National de la Recherche Scientifique, and later to Brazil, where he was hired as a professor at the University of Recife.
@@
  • Francisco Paulo Pinto Leite (1902-2000), who was a lecturer at the Sorbonne from 1931 to 1933. One of the founders of the Portuguese Mathematical Society, he was appointed Full Professor at ISCEF in 1940, retiring in 1973. There are various versions of the course notes for the Statistics course he taught at ISCEF, and he was also the first lecturer, in 1952-53, of the chair of Econometrics, with Henri Guitton (University of Dijon) and José de Castañeda (University of Madrid), having as assistant Manuel Jacinto Nunes (further information in Machado and Santos Silva, 2002 [39]).
@@
    It should be noted that 1953 saw the publication of Trabalhos do Seminário de Econometria Dirigido pelo Prof. H. O. Wold and the translation of Wold’s book on dynamic systems by Jacinto Nunes, and that Mira Fernandes (who prefaced the book by Mário Santos and Zaluar Nunes) consistently supported research in Statistics and Econometrics, see the series of eight volumes of Estudos de Matemática, Estatística e Econometria [Studies in Mathematics, Statistics and Econometrics] that he edited between 1956 and 1964.
@@
  • José Sebastião e Silva (1914-1972), a fellow of the Instituto de Alta Cultura in Rome from 1942 to 1946, in 1951 won the public competition for the position of Full Professor at the Instituto Superior de Agronomia, taking charge of the chairs of General Mathematics and Calculus of Probabilities. For the latter chair, in 1954-55, he prepared didactic texts, collected in the course notes Cálculo das Probabilidades [Calculus of Probabilities] published at ISA, with an appendix from 1957-58 on correlation and regression, adjustment by the method of least squares, and Student’s $t$ and Fisher-Snedecor’s $F$ statistics, cf. Neves (2005) [40], pp.222–223. In 1991 Calouste Gulbenkian Foundation republished this work, now unfortunately out of print.
@@
Perhaps as a result of the influence of the work developed by Sir Ronald Fisher and his collaborators at Rothamsted, new areas of Statistics were developed at ISA, particularly in the area of Experimental Design (Costa, 1933; Carvalho, 1945, 1946) and Linear Models (Pestana, 1957).
@@
Professor Varennes e Mendonça taught Probabilities at ISA before the appointment of Professor Sebastião e Silva, publishing a brief course on Noções de Cálculo das Probabilidades [Notions of Probability Calculus] (1950), for details cf. Neves (2005) [40], pp. 224–225. Varennes e Mendonça early on realized the difficulties of translating statistical terms, publishing an interesting article on terminology in 1942.
@@
  • Murteira and Madureira (1958) supplied interesting translations, but it wasn’t until much later that the Specialised Committee on Statistical Nomenclature of the Portuguese Statistical Society and the Brazilian Statistical Association, whose members were Carlos Daniel Paulino (President); Dinis Pestana; João Branco; Lucília Carvalho; Alfredo Egídio dos Reis; Júlio Singer (Chairman); Lúcia Barroso; Wilton Bussab, prepared the Statistics Glossary English-Portuguese, Portuguese Statistical Society / Brazilian Statistical Association https://spestatistica.pt/glossario, which continues to evolve; see also Ventura, C.; Gomes, M.I.; Pestana, D.; Pestana, P. (2013). Glossary of Statistical Terms: German, French, English, Portuguese, INE, Lisbon.
@@
The translation, in 1945, of Snedecor’s Statistical Methods Applied to Experiments in Agriculture and Biology certainly had a major impact on the development of Statistics in Portugal. The popularisation of interest in Probability and Statistics, with an increase in the number of students taking courses in these areas, finally aroused the interest of publishers.
@@
  • Publicações Europa-América pioneered this move to publish statistical books, with the publication of Gustavo de Castro’s books in the collection it Saber [Knowledge] (1959, 1965), an author chosen probably because he had previously published a book on Inferência Estatística [Statistical Inference] at LNEC [National Laboratory of Civil Engineering] in 1953.
@@
This was followed by the publication of the translation of Tippett’s Statistics (1955), and in 1973-75 Europa-América published the two volumes of P. Dagnelie’s Théorie et Méthodes Statistiques: Applications Agronomiques, translated by A. St. Aubyn, and in 1982 the translation of M. L. Lévy’s Comprendre les Statistiques.
@@
  • Editorial Estúdios COR also ventured out with the publication in 1968 of a translation of Gnedenko and Khinchine’s book An Elementary Introduction to the Theory of Probability, with a preface and translation by António Simões Neto; and Despertar published a translation of one of the highest quality and most successful introductory books, Moroney’s Facts from Figures, in 1969.
@@
The books by Guimarães (1904) and Sousa Pinto (1913) were analysed by Rui Santos, and those by Varennes e Mendonça (1950), Sebastião e Silva (1954) and Pestana (1957) by Manuela Neves [40], as already mentioned.
Of the others, we think the publications by Santos and Zaluar Nunes (1933), Braumann (1958), Reis and Sarmento (1960) and Dias Agudo (1961) deserve special mention.
The publications by E. Tamagnini (1937) and W. L. Stevens (1942-45) in the Revista da Faculdade de Ciências da Universidade de Coimbra [Review of the Faculty of Sciences of Coimbra University] also deserve special mention.
  • The book by Santos and Zaluar Nunes (1933) has a modernity that contrasts sharply with the 1932 book by Bueno y Martins, which has a perspective closer to official statistics.
@@
    • However, when it comes to descriptive statistics, Bueno y Martins’ book is more detailed, for example, in addition to the arithmetic mean, he considers geometric, harmonic and weighted averages; and in what regards graphical representations, Bueno y Martins discusses cartograms, stereograms, presenting various interesting pictograms, and discusses as well as association and contingency, and correlation.
@@
    It should be noted, however, that some of the more elaborate graphics in Bueno y Martins (1932) come from other sources. For example, the extra text between pages 138 and 139 is a reproduction of the cover of the August 1930 issue of the Boletim da Direção Geral de Estatística [Bulletin of the General Directorate of Statistics], see the Catálogo Bibliográfico [Bibliographic Catalogue], p.19, published by INE in 2013. Bueno y Martins’ main statistical reference is the monograph by Yule (the bibliographical sources used are indicated at the end of each section, which is very useful), and it is worth noting that Bueno y Martins indicates several books on Probability in the final bibliography, including A Treatise on Probability by Maynard Keynes.
@@
Brilhante, M. F.; Pestana, D.; Rocha, J.; Rocha, M. L.; Velosa, S. (2001) 2ª ed. 2011, Inferência Estatística sobre a Localização Usando a Escala [Statistical Inference about Location Using Scale]
@@
Brilhante, M. F. (2017) An Introduction to Meta-Analysis [An Introduction to Meta-Analysis]
Amado, C.; Nunes, C.; Sardinha, A. (2019) Análise Estatística de Dados Financeiros [Statistical Analysis of Financial Data]
@@
Actas I Colóquio de Estatística e Investigação Operacional, Lisboa, Centro de Estatística e Aplicações, 1977.
Actas II Colóquio Estatística e Investigação Operacional, Fundão/Covilhã, CEAUL/SPEIO, 1981.
Actas III Colóquio Estatística e Investigação Operacional, CEAUL/SPEIO, Tróia, 1985.
@@
Afirmar a Estatística: um Desafio para o Século XXI [Affirming Statistics: a Challenge for the 21st Century], Atas do VI Congresso da SPE, 1999, Tomar; ed.  Carlos Daniel Paulino, António Pacheco, Ana Pires e Ferreira da Cunha.
Um Olhar sobre a Estatística [A Look at Statistics], Actas do VII Congresso da SPE, 2001, Ofir; ed.  Pedro Oliveira e Emília Athayde.
@@
Estatística Jubilar [Jubilee Statistics], Actas do XII Congresso Anual da SPE, 2005, Évora; ed.  Carlos Braumann, Paulo Infante, Manuela M. Oliveira, Russell Alpízar-Jara e Fernando Rosado.
@@
Estatística, Ciência Interdisciplinar [Statistics, Interdisciplinary Science], Actas do XIV Congresso da SPE, 2007, Covilhã; ed.  Maria Eugénia Ferrão, Célia Nunes e Carlos Braumann.
@@
Estatística: Arte de Explicar o Acaso [Statistics: Art of Explaining Chance], Actas do XVI Congresso Anual da SPE, 2009, Vila Real; ed. Irene Oliveira, Elisabete Correia, Fátima Ferreira, Sandra Dias e Carlos Braumann.
@@
New Frontiers in Statistics and Data Science, Springer Proceedings in Mathematics & Statistics, Springer, XXVI Congresso da SPE, 2024, Guimarães (ed. Lígia Henriques-Rodrigues, Raquel Menezes, Luís Meira Machado, Miguel de Carvalho e Susana Faria), in print.
@@
Index of the Manual de Estatística Médica (Manual of Medical Statistics) by Carlos Santos Reis and Alexandre Sarmento.
@@
   5 – Comparison of main position measurements 142

### v2

@@ 1 Introduction
The teaching of Statistics associated to Political Economy persisted at the Faculty of Law of the University of Coimbra, giving rise to the publication of various texts (Gouveia et al. [67], Marnoco e Souza999  The incomplete copy from FDUL did not allow us to access some parts of the book; it should be noted that it has a section on “the sceptics of Statistics”, a subject that is currently widely discussed. In Chapter III (Statistical Laws), sections 59 and 60 are respectively “Law of large numbers” and “Law of small numbers”. Poisson, wishing that his result on the convergence of a succession of binomials to Poisson to be known as the law of small numbers, renamed “law of large numbers” Jacques Bernoulli’s “golden theorem” on the convergence of arithmetic means towards the expectation. Only the expression “law of large numbers” persisted. [101], Silva and Rocha [162], Rocha and Monteiro [145], Pereira [135]).
On the other hand, at the Anthropology Institute of the University of Coimbra, Eugénio Tamagnini had an innovative perspective on the contribution of Statistics in Biometrics research. In Revista da Faculdade de Ciências da Universidade de Coimbra, Tamagnini [181] published an innovative work in Portugal on analysis of variance. Tamagnini also hired W. L. Stevens, a disciple of Ronald Fisher. Xavier da Cunha (1992) [208] comments that “Tamagnini’s interest in the statistical treatment of data obtained from various populations and in comparing them led him to hire a foreign professor of statistics, Professor W. L. Stevens, who collaborated with him on blood groups and also organised free courses in statistics for students of the exact and natural sciences. Stevens had been a brilliant disciple and collaborator of Dr. Fisher, having stayed in Coimbra ‘making us familiar with statistical methods and their most convenient application’, in Tamagnini’s words. This initiative was undoubtedly very useful.
@@
Fonseca [50] states that “It is in this context of the necessity to improve the statistical treatment of anthropometric, physiological and demographic data, that one understands the arrival of W. L. Stevens, assigned ‘as a statistician, to guide the application work of modern statistical methods to Biological Sciences and organise initiation courses for professors and students of the same methods’. The ‘Elementary Course of Modern Statistical Methods Applicable to Scientific Investigation’, headed by professor W. L. Stevens, functioned between 1942 and 1944, and is linked to a set of published studies in the series ‘Questões de Método’ [‘Methodogical Issues’]. […] As for the years W. L. Stevens spent in Coimbra, the data which allows us to inquire about his teaching activity is frankly scarce. We know that the course he led functioned between 1942 and 1944, but the apparent absence of Stevens’ curriculum vitae in the archives of the University of Coimbra, and, moreover, the inexistence of annuals of the same institution for the period between 1943 and 1947 (encompassing, thus, the years during which the course led by W. L. Stevens was lectured), conditioned our investigation. What was the programme of the course? Who attended it? What bibliography was recommended? Why was it extinct after only two years? These are just some of the questions that (at least) for now, will remain without an answer.”.
Between 1942 and 1945 Stevens published a remarkable set of works which he categorised as Questions of Method — namely on the $\chi^{2}$ significance test, discriminant analysis, statistical distributions, parameters estimation —, disseminating innovations in mathematical statistics by Karl Pearson, ‘Student’, Ronald Fisher, Egon Pearson, Jerszy Neyman and their schools.
Stevens also published, with Aureliano Quintanilha and Hugo Ribeiro [141] an interesting article on the application of Probability Calculus in the investigation of non sexual reproduction by plasmogamy followed by karyogamy in the higher fungi basidiomycetes, in the Gazeta de Matemática (both in the index and in the article the initials W and L are interchanged).
@@
  • Manuel Zaluar Nunes (1907-1967), who started as an assistant at ISCEF, before going to study in Paris as a fellow of the Instituto de Alta Cultura, published in 1933 with Mário Santos a concise exposition of Elementos de Cálculo das Probabilidades e de Estatística Matemática [149] [Elements of Probability Calculus and Mathematical Statistics]: written in accordance with the program of the 2nd chair of the Instituto Superior de Ciências Económicas e Financeiras. Founder of the Gazeta de Matemática, of Portugaliae Mathematica, and of the Sociedade Portuguesa de Matemática [Portuguese Mathematical Society], he was an Assistant Professor at FCUL and later a Full Professor at ISA, a position from which he was dismissed in 1947 for political reasons, having to migrate first to France, where he was a research assistant at the Centre National de la Recherche Scientifique, and later to Brazil, where he was hired as a professor at the University of Recife.
@@
  • Francisco Paulo Pinto Leite (1902-2000) was lecturer at Sorbonne from 1931 to 1933. One of the founders of the Portuguese Mathematical Society, he was appointed Full Professor at ISCEF in 1940, retiring in 1973. There are various versions of the course notes for the Statistics course he taught at ISCEF, and he was also the first lecturer, in 1952-53, of the chair of Econometrics, with Henri Guitton (University of Dijon) and José de Castañeda (University of Madrid), having as assistant Manuel Jacinto Nunes (further information in Machado and Santos Silva [97]).
@@
    It should be noted that 1953 saw the publication of Trabalhos do Seminário de Econometria Dirigido pelo Prof. H. O. Wold [206] [Works presented at the Seminar of Econometry drected by Professor H. O. Wold] and the translation of Wold’s book on dynamic systems [206] by Jacinto Nunes, and that Mira Fernandes (who prefaced the book by Mário Santos and Zaluar Nunes) consistently supported research in Statistics and Econometrics, see the series of eight volumes of Estudos de Matemática, Estatística e Econometria [Studies in Mathematics, Statistics and Econometrics] [108] that he edited between 1956 and 1964.
@@
  • José Sebastião e Silva (1914-1972), a fellow of the Instituto de Alta Cultura in Rome from 1942 to 1946, in 1951 won the public competition for the position of Full Professor at the Instituto Superior de Agronomia, taking charge of the chairs of General Mathematics and of Calculus of Probabilities. In 1954-55, he prepared didactic texts, collected in the course notes Cálculo das Probabilidades[158] [Calculus of Probabilities] published at ISA, with an appendix from 1957-58 on correlation and regression, adjustment by the method of least squares, and Student’s $t$ and Fisher-Snedecor’s $F$ statistics, cf. Neves [125], pp.222–223. In 2000 Calouste Gulbenkian Foundation republished this work [159], now unfortunately out of print.
@@
Perhaps as a result of the influence of the work developed by Sir Ronald Fisher and his collaborators at Rothamsted, new areas of Statistics were developed at ISA, particularly in the area of Experimental Design (Costa [36], Carvalho [28, 29]) and Linear Models (Pestana [137]).
@@
  • Professor Varennes e Mendonça, who graduated in Agronomy Engineering at ISA, taught Probabilities at ISA before the appointment of Professor Sebastião e Silva, publishing a brief course on Noções de Cálculo das Probabilidades [194] [Notions of Probability Calculus], for details cf. Neves [125], pp. 224–225. Varennes e Mendonça early on realized the difficulties of translating statistical terms, publishing an interesting article [193] on terminology in 1942.
@@
  • Murteira and Madureira [122] supplied interesting translations, but it wasn’t until much later that the Specialised Committee on Statistical Nomenclature of the Portuguese Statistical Society and the Brazilian Statistical Associationprepared the Statistics Glossary English-Portuguese [133], https://spestatistica.pt/glossario, which is periodically updated. See also Ventura [197] et al’s [Glossary of Statistical Terms: German, French, English, Portuguese].
@@
The translation, in 1945, of Snedecor’s Statistical Methods Applied to Experiments in Agriculture and Biology [163] certainly had a major impact on the development of Statistics in Portugal. The popularisation of interest in Probability and Statistics, with an increase in the number of students taking courses in these areas, finally aroused the interest of publishers.
@@
  • Publicações Europa-América pioneered this move to publish statistical books, with the publication of Gustavo de Castro’s books [32, 33] in the collection it Saber [Knowledge] (1959, 1965), an author chosen probably because he had previously published a book on Inferência Estatística [31] [Statistical Inference] at LNEC [National Laboratory of Civil Engineering] in 1953.
@@
This was followed by the publication of the translation of Tippett’s Statistics [190] in 1955 (reprinted in 1968), and in 1973-75 Europa-América published the two volumes of P. Dagnelie’s Théorie et Méthodes Statistiques: Applications Agronomiques [37], translated by A. St. Aubyn, and in 1982 the translation of M. L. Lévy’s Comprendre les Statistiques [96].
@@
  • Editorial Estúdios COR also ventured out with the publication in 1968 of Gnedenko and Khinchine’s book Elementary Introduction to the Theory of Probability [58], with preface and translation by A. Simões Neto. Despertar published a translation of one of the highest quality and most successful introductory books, Moroney’s Facts from Figures [113], in 1969.
@@
The books by Guimarães [75] and Sousa Pinto [173] were analysed by Rui Santos, and those by Varennes e Mendonça [194], Sebastião e Silva [158] and Pestana [137] by Manuela Neves [125], as already mentioned.
Of the others, we think the publications by Santos and Zaluar Nunes [149], Braumann [18], Reis and Sarmento [143] and Dias Agudo [42] deserve special mention.
The publications by E. Tamagnini [181] and W. L. Stevens [174, 175, 176, 177, 178] in the Revista da Faculdade de Ciências da Universidade de Coimbra [Review of the Faculty of Sciences of Coimbra University] also deserve special mention.
  • The 1933 book by Santos and Zaluar Nunes [149] has a modernity that contrasts sharply with the 1932 book by Bueno y Martins [23], which has a perspective closer to official statistics.
@@
    • However, when it comes to descriptive statistics, Bueno y Martins’ book is more detailed, for example, in addition to the arithmetic mean, he considers geometric, harmonic and weighted averages.
@@
    In what regards graphical representations, Bueno y Martins discusses cartograms, stereograms, presenting various interesting pictograms, and discusses as well as association and contingency, and correlation. It should be noted, however, that some of the more elaborate graphics in Bueno y Martins [23] come from other sources. For example, the extra text between pages 138 and 139 is a reproduction of the cover of the August 1930 issue of the Boletim da Direção Geral de Estatística [Bulletin of the General Direction of Statistics], see the Catálogo Bibliográfico [Bibliographic Catalogue], p.19, published by INE in 2013 [82]. Bueno y Martins’ main statistical reference is the monograph by Yule (the bibliographical sources used are indicated at the end of each section, which is very useful), and it is worth noting that Bueno y Martins indicates several books on Probability in the final bibliography, including A Treatise on Probability by Maynard Keynes.
@@
Brilhante, M. F.; Pestana, D.; Rocha, J.; Rocha, M. L.; Velosa, S. (2001) 2ª ed. 2011, Inferência Estatística sobre a Localização Usando a Escala [Statistical Inference about Location Using the Scale]
@@
Brilhante, M. F. (2017) Uma Introdução à Meta-Análise [An Introduction to Meta-Analysis]
Amado, C.; Nunes, C.; Sardinha, A. (2019) Análise Estatística de Dados Financeiros [Statistical Analysis of Finance Data]
@@
Actas, I Colóquio de Estatística e Investigação Operacional, Lisboa, Centro de Estatística e Aplicações, 1977.
Actas, II Colóquio Estatística e Investigação Operacional, Fundão/Covilhã, CEAUL/ SPEIO, 1981.
Actas, III Colóquio Estatística e Investigação Operacional, CEAUL/SPEIO, Tróia, 1985.
@@
Afirmar a Estatística: um Desafio para o Século XXI [Asserting Statistics: a Challenge for the 21st Century], Atas do VI Congresso da SPE, 1999, Tomar; ed.  Carlos Daniel Paulino, António Pacheco, Ana Pires e Ferreira da Cunha.
Um Olhar sobre a Estatística [A Look upon Statistics], Actas do VII Congresso da SPE, 2001, Ofir; ed.  Pedro Oliveira e Emília Athayde.
@@
Estatística Jubilar [Jubilar Statistics], Actas do XII Congresso Anual da SPE, 2005, Évora; ed.  Carlos Braumann, Paulo Infante, Manuela M. Oliveira, Russell Alpízar-Jara e Fernando Rosado.
@@
Estatística, Ciência Interdisciplinar [Statistics, an Interdisciplinary Science], Actas do XIV Congresso da SPE, 2007, Covilhã; ed.  Maria Eugénia Ferrão, Célia Nunes e Carlos Braumann.
@@
Estatística: Arte de Explicar o Acaso [Statistics: the Art of Explaining Chance], Actas do XVI Congresso Anual da SPE, 2009, Vila Real; ed. Irene Oliveira, Elisabete Correia, Fátima Ferreira, Sandra Dias e Carlos Braumann.
@@
New Frontiers in Statistics and Data Science, Springer Proceedings in Mathematics & Statistics, Springer, XXVI Congresso da SPE, 2024, Guimarães (ed. Lígia Henriques-Rodrigues, Raquel Menezes, Luís Meira Machado, Miguel de Carvalho e Susana Faria).
@@
Table of Contents of the Manual de Estatística Médica [143] [Manual of Medical Statistics] by Carlos Santos Reis and Alexandre Sarmento.
@@
5 – Comparison of the main location measurements 142